Hello, what is the x-intercept of the problem? -5x+ y=3
Answer by Alan3354(69443)

what is the x-intercept of the problem? -5x+ y=3
-----------------
It's where y = 0
-5x + 0 = 3
x = -3/5, or the point (-3/5,0)